OpenStudy (anonymous):

please help! what is the domain and range for x+2= square root of y?

OpenStudy (anonymous):

domain is all real numbers range i all numbers greater than 0

OpenStudy (anonymous):

thank you. can you explain?

OpenStudy (anonymous):

so, sqrt(y) = x+2 is the same as y = (x + 2)^2

OpenStudy (anonymous):

there are no numbers that cant be pluged into x, so the domain is all real numbers

OpenStudy (anonymous):

but, you will only get positive numbers back, so the range is all real numbers greater than 0

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

np